# Excitable quantum systems: the bosonic avalanche laser

### v2

@@ Abstract
Acknowledgments.—We thank Mathias Marconi, Lukas Schamriss and Kirill Fedorov for helpful discussions. This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
Supplementary Material for “Excitable quantum systems: the bosonic avalanche laser”
II ASIP dynamics and population profile
@@
In the limit of large $\kappa_{c}$, the cavity amplitude can be set to zero, and the dynamics described by (3) becomes a process for the ladder of modes only, described by the ASIP model Garbe et al. (2024). Here we recall a few of its main properties. The ASIP dynamics takes the form of a conservation law, with a mean-field current $J_{p,p+1}=\Gamma n_{p}(1+n_{p+1})$ between modes $p$ and $p+1$. When the population distribution is smooth, we may approximate the discrete populations $n_{p}$ by a continuous field $n(x)$. The dynamics then becomes a continuous partial differential equation:
$\partial_{t}n=\Gamma\left[(1+2n)\partial_{x}n+\frac{1}{2}\partial^{2}_{x}n\right]$
@@
also known as Burgers’ equation [ref?]. This dynamics is characterized by a competition between non-linear advective transport and ordinary diffusion. In the transient regime, this leads to well-known non-linear transport effects, such as shock waves and fronts. However, in the presence of boundary drive and dissipation, and in the steady-state regime, this competition generates an alternating pattern, whereby the population shows oscillations around the population of the first site $n_{1}$, with an amplitude growing near-exponentially as one goes towards the last site. This pattern of staggered population and accumulation on the ladder boundary, which we described as bosonic skin effect Garbe et al. (2024), is a consequence of the non-linearity of the transport, induced by the bosonic statistics of the particles.
@@
More quantitatively, one may rewrite the steady-state condition as $j_{p,p+1}=J\hskip 5.0pt\forall p$, ie, the steady-state imposes an homogeneous, constant current. This current can be obtained from the boundary condition, and will depend on the specific pumping mechanism; in the case of infinite-temperature reservoir $\zeta=1/2$ which we consider in the main text, we get simply $J=\gamma_{\textsl{g}}$. The steady-state condition then becomes a recurrence relation for the population: Quantitatively, the population of the first site can be calculated from the left boundary condition: $n_{p}=\frac{J/\Gamma}{1+n_{p+1}}$, which gives the alternating pattern we observe. The population on the first and last sites of the chain can be obtained through the boundary condition:
$\gamma_{\textsl{g}}=\Gamma n_{1}(1+n_{1})$ $\Rightarrow n_{1}\sim\sqrt{\frac{\gamma_{\textsl{g}}}{\Gamma}}$
$\gamma_{\textsl{g}}=\kappa_{\ell}n_{N}$ $\Rightarrow n_{N}=\frac{\gamma_{\textsl{g}}}{\kappa_{\ell}}$
Here, we have assumed that the chain is long enough for the oscillations to be negligible near the first site, which allows to rewrite the current as $\Gamma n_{1}(1+n_{2})\sim\Gamma n_{1}(1+n_{1})$. Note that the "parity" of the oscillations, ie, the sign of the staggered population $n_{\text{stag}}=\sum_{p}(-1)^{p}(n_{p}-n_{1})$, will be simply given by the sign of $n_{N}-n_{1}$. Note that this condition is not related to the existence of a pulsing phase; by tuning the loss rate $\kappa_{\ell}$ on the last ladder mode, one may change the parity of the oscillations while maintaining an empty cavity.
@@
$c=\Gamma(1+2n_{1})\sim 2\sqrt{\Gamma\gamma_{\textsl{g}}}$
@@
The presence of the $n_{1}$ factor indicates that the effective speed is renormalized by the average filling of the ladder, which is again a consequence of the stimulation of the transport by other bosons in the ladder. This quantity provides the typical time-scale at which perturbations propagate along the ladder, and eventually subside; in the pulsing phase, it provides the scale to compare the oscillation period to (see next section).
@@
In the lasing phase, the cavity population will settle to a non-zero value. The dynamics in the chain will once more be described by the ASIP model, with an effective hopping rate $\Gamma(1+\lvert\alpha_{c}\rvert^{2})$ rescaled by the cavity population. The steady-state will once more display a modified staggered pattern.
III Details on pulsing dynamics and scaling of the period
@@
In this section, we elaborate on the dynamics of the ladder in the pulsing phase, and the determination of the oscillation period. In Fig.7, we show snapshots of the cavity population distribution at different times, for two different values of the cavity loss rate. These corresponds to points in the pulsing phase close to the boundaries with the empty and lasing phases, respectively. In the latter case (bottom plot), we observe the kind of mechanism we described in the main text: starting from a mostly empty ladder (the rightmost site population that can be observed in the plot is a leftover from the previous pulse), we observe a packet slowly forming and propagating on the first few modes of the ladder. As described in the main text and the previous section, the speed of motion of this packet can be estimated as $c\sim\sqrt{\gamma_{\textsl{g}}\Gamma}$. As this packet propagates along the ladder, an increased number of modes get populated and engage in the hopping dynamics. This, in turns mean a larger pumping for the cavity; once the total pumping rate exceeds the loss rate $\kappa_{c}$, one observes a burst-like increase in $n_{c}$ (right-side plot). This, in turns, triggers a sudden acceleration of the packet, which quickly moves all the way through the ladder to the final site.
@@
For high values of $\kappa_{c}$ (top plot in Fig.7), we observe a similar, but slightly different process: instead of starting from an empty ladder, we have a non-zero filling, such that the total cavity pumping already exceeds the threshold value $\kappa_{c}$. We observe, therefore, a rapid increase in the cavity population, resulting once again in an acceleration of the ladder dynamics. Contrary to the previous case, however, the resulting effect is now a dip in the ladder population on the left-hand side. To understand this, consider the dynamics of the leftmost site; in the initial configuration we considered, we have a local equilibrium between the incoming drive current $\gamma_{\textsl{g}}$ and the current $\Gamma(1+\lvert\alpha_{c}\rvert^{2})n_{1}(1+n_{2})$ flowing to the second site. When the cavity population spikes, the latter increase, but not the former, resulting in a population loss on the first site (modes in the center of the ladder, by contrast, are initially unaffected, since both their incoming and outgoing currents are amplified). This initial quench on the first site then quickly affect the neighboring modes, creating the depleted region we observe on the first few modes of the ladder. The decrease in population leads to a lower overall current, making the cavity population decay once more. We are therefore left with a depleted region in the ladder, which is going to propagate at the same effective speed $c$ we discussed above. The depleted region will propagate until it reaches the right edge; once it starts subsiding, the population moves back towards its initial constant filling configuration, and the current starts increasing again, restarting the cycle.
@@
To summarize, these two opposite regimes of high and low $\kappa_{c}$ exhibit behaviors that are more or less mirror images of one another: in the former, we have a slow creation of a population wave, followed by rapid propagation through the ladder; in the latter , we have a rapid creation of a depletion wave, followed by slow propagation through the ladder. The slow propagation of both kind of waves, however, occurs with the same typical speed $c\sim\sqrt{\Gamma\gamma_{\textsl{g}}}$. Since each oscillation period involves both slow and fast processes, the overall duration of the period will be governed by the former.
In the case of high $\kappa_{c}$, the duration of this slow protocol can be estimated by the argument given in the main text: the number of occupied modes grows like $ct$, giving a total current $J_{cum}=c\gamma_{\textsl{g}}t$; the pulses are triggered when the total current becomes of the order of $\kappa_{c}$, ie, at time $c\tau=\kappa_{c}/\gamma_{\textsl{g}}$. Hence, this argument suggests that the period, rescaled by the effective speed $c\sim\sqrt{\gamma_{\textsl{g}}\Gamma}$, should be compared against the rescaled gain $\gamma_{\textsl{g}}/\kappa_{c}$. In Fig.8, we present a fuller view of these results, displaying the unrescaled and rescaled periods for multiple values of the hopping rate, cavity loss rates, and gain. We observe a very good collapse of all curves indicating that the ansatz $c\tau=f(\kappa_{c}/\gamma_{\textsl{g}})$ works quite well for both high and low $\kappa_{c}$. However, a more detailed analysis of the function $f$ reveals that it deviates from the behavior $f(\kappa_{c}/\gamma_{\textsl{g}})=\kappa_{c}/\gamma_{\textsl{g}}$ that one would expect from the previous argument. The precise behavior of this function is not fully clear and will be investigated in future works.
Figure 8: Left: unrescaled period versus gain rate, setting $N=10$ and $\kappa_{\ell}$ constant, and taking multiple values for the hopping rate and cavity loss: $\Gamma/\kappa_{\ell}=0.5,0.1,0.05$ (blue, green, and red symbols, respectively), and $\kappa_{c}/\kappa_{\ell}=1,5,10,25,50$(squares, circles, crosses, stars, and diamonds, respectively). Right: rescaled period versus rescaled pumping, with same parameters. We used this plot for the inset of Fig.3 in the main text.
IV Monte-Carlo simulations
In the absence of any additional Hamiltonian terms, the master equation in Eq. (1) is diagonal in the Fock basis $|\{\vec{n}\}\rangle=|n_{c};n_{1},n_{2}..,n_{N}\rangle$,where $n_{c}$ and $n_{1,..N}$ denote the population of the cavity and the modes in the ladder (i.e., the gain medium), respectively. Therefore, the diagonal elements of the density operator, $P(\{\vec{n}\},t)=\langle\{\vec{n}\}|\hat{\rho}(t)|\{\vec{n}\}\rangle$, which describe the probabilities of different particle configurations, evolve independently from the off-diagonal elements, according to the following equation:
@@ IV Monte-Carlo simulations
Where the $dH_{q}$, $dG_{\lambda}$, and $dL_{\lambda}$ are independent random variables, taking binary values $\{0,1\}$, and indicating that hopping, gain, and loss event, respectively, took place. In a short enough time interval $dt$, the respective probabilities that these variables assume value $1$ are $p(dH_{q}=1)=\Gamma(1+n_{c})n_{q+1}(1+n_{q})dt$, $p(dG_{1}=1)=\gamma_{\textsl{g}}(1+n_{1})dt$, and $p(dL_{\lambda}=1)=\kappa_{\lambda}n_{\lambda}dt$.
The simulation is then performed using standard Gillespie algorithm Gillespie (1977). By starting from a given initial configuration, $\{n_{p}(t=0)\}$, and evolving a total number of $\mathcal{N}_{t}$ stochastic trajectories in time, we can approximate the expectation value of any function of operators $\hat{n}_{p}$ by an ensemble average. For example,
@@
V Losses in the ladder
An implementation of the model described here will necessarily involve dissipation channels beyond the one we described. As we discussed in the End Matter, channels inducing only loss of coherence will not affect the population dynamics, and in particular, will have no effect on the pulsing behavior. The main detrimental process, therefore, is the loss of particles within the ladder of modes. For an implementation in superconducting circuits, with individual modes are realized using high-$Q$ microwave cavities, such losses can be as low as $\kappa_{0}=10\text{kHz}$ Frunzio et al. (2005); Reagor et al. (2013); Somoroff et al. (2023). Given the parameters we have for the implementation, we can thus expect $\kappa_{0}\sim 0.1-1\Gamma$. We performed mean-field simulations taking such process into account; in Fig.9, we display the phase diagram we obtain, using the same parameters as in Fig.2 in the main text, and $\kappa_{0}/\Gamma$ ranging from $0.1$ to $10$. We see that the phase diagram is almost unaffected for all but the largest values of $\kappa_{0}$, and that, even in this case, we can clearly distinguish the three phases we discussed. Based on this mean-field analysis, we expect that none of our findings will be critically affected by the typical levels of losses one could expect in superconducting circuits.
VI Coherence resonance with other pumping mechanisms
@@
In the main paper, we have mostly focused on a pumping mechanism with $\zeta=1/2$, which corresponds to equal gain and loss on the first ladder mode. At the level of the amplitude of the mode, this would correspond to driving by a random force. The phenomenon of coherence resonance, however, can also be achieved with other pumping mechanisms. In Fig.10, we show the trajectories, spectra, and coherence factor obtained with $\zeta=1$, which corresponds to pure gain in the first ladder site (or, at the level of the mode amplitude, a stochastic force with non-zero average). The same basic features can be observed: as the gain rate is increased, the cavity population shows an increasing number of burst, forming a quasi-regular behavior, before entering a lasing regime. The noise spectra show the same basic behavior as the ones displayed in Fig.4 of the main text, and the plot of $\beta$ against the gain rate again shows the maximum indicative of coherence resonance.

### v3

@@ Acknowledgements.
We thank Mathias Marconi, Lukas Schamriss and Kirill Fedorov for helpful discussions. This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
Appendix A ASIP dynamics and population profile
@@
In the limit of large $\kappa_{c}$, the cavity amplitude $\alpha_{c}$ can be set to zero, and the dynamics described by (4) becomes a process for the ladder of modes only, described by the so-called ASIP model [7]. When the population distribution is smooth, we may approximate the discrete populations $n_{p}$ by a continuous field $n(x)$, that obeys the partial differential equation:
$\partial_{t}n=\Gamma\left[(1+2n)\partial_{x}n+\frac{1}{2}\partial^{2}_{x}n\right]$ (12)
@@
This dynamics is characterized by a competition between ordinary diffusion and non-linear advective transport, induced by the bosonic statistics of the particles. In the presence of boundary drive and dissipation, and in the steady-state regime, this competition generates a pattern of oscillating population, whose amplitude grows near-exponentially as one goes towards the last site, and which we dubbed bosonic skin effect [7].
@@
Quantitatively, one may rewrite the steady-state condition as $J_{p,p+1}=J_{ss}\hskip 5.0pt\forall p$, i.e., the steady-state imposes a homogeneous, constant current, which can be obtained from the boundary condition. In the case $\zeta=1/2$, which we consider in the main text, we get simply $J_{ss}=\gamma_{\textsl{g}}$. The steady-state condition then becomes a recurrence relation for the population: $n_{p}=\frac{J/\Gamma}{1+n_{p+1}}$, from which we can derive the alternating pattern. Near the left boundary, in which the oscillation amplitude vanishes, we have $n_{2}\sim n_{1}$, from which we can infer $n_{1}\sim\sqrt{\gamma_{\textsl{g}}/\Gamma}$. The population on the last site can also be obtained through the boundary condition, giving us access to the full oscillation profile (see [7] for the full analytical expression).
@@
$c=\Gamma(1+2n_{1})\sim 2\sqrt{\Gamma\gamma_{\textsl{g}}}.$ (13)
@@
The presence of the $n_{1}$ factor indicates that the effective speed is renormalized by the average filling of the ladder, which is again a consequence of the stimulation of the transport by other bosons in the ladder. This quantity provides the typical time-scale at which perturbations propagate along the ladder and eventually subside; in the pulsing phase, it provides the scale to compare the oscillation period to (see section C).
@@
Finally, in the lasing phase, the cavity population will settle to a non-zero value. The dynamics in the chain will once more be described by the ASIP model, with an effective hopping rate $\Gamma(1+\lvert\alpha_{c}\rvert^{2})$ rescaled by the cavity population. The steady-state will once more display a modified staggered pattern determined by this effective hopping rate.
Appendix B Losses in the ladder
An implementation of the model described here will necessarily involve dissipation channels beyond the one we described. As we discussed in the End Matter, channels inducing only loss of coherence will not affect the population dynamics, and in particular, will have no effect on the pulsing behavior. The main detrimental process, therefore, is the loss of particles within the ladder of modes. For an implementation in superconducting circuits, in which individual modes are realized using high-$Q$ microwave cavities, such losses can be as low as $\kappa_{0}=10\text{kHz}$ [5, 28, 31]. Given the parameters we have for the implementation, we can thus expect $\kappa_{0}\sim 0.1-1\Gamma$. We performed mean-field simulations taking such processes into account; in Fig.8, we display the phase diagram we obtain, using the same parameters as in Fig.2 in the main text, and $\kappa_{0}/\Gamma$ ranging from $0.1$ to $10$. We see that the phase diagram is almost unaffected for all but the largest values of $\kappa_{0}$, and that, even in this case, we can clearly distinguish the three phases we discussed. Based on this mean-field analysis, we expect that none of our findings will be critically affected by the typical levels of losses one can expect in superconducting circuits.
Appendix C Details on pulsing dynamics and scaling of the period
@@
In this section, we elaborate on the dynamics of the ladder in the pulsing phase, and on the determination of the oscillation period. In Fig.9, we show snapshots of the cavity population distribution at different times, for two different values of the cavity loss rate. These correspond to points in the pulsing phase close to the boundaries with the empty and lasing phases, respectively. In the latter case (bottom plot), we observe the kind of mechanism we described in the main text: starting from a mostly empty ladder (the rightmost site population that can be observed in the plot is a leftover from the previous pulse), we observe a packet slowly forming and propagating on the first few modes of the ladder. As described in the main text and in Appendix A, the speed of motion of this packet can be estimated as $c\sim\sqrt{\gamma_{\textsl{g}}\Gamma}$. If we assume the dynamics within the packet itself to be quasi-stationary, the current between each pair of sites is given by the steady-state condition $J_{p,p+1}=J_{ss}=\gamma_{\textsl{g}}$. As this packet propagates along the ladder, an increased number of modes get populated and engage in the hopping dynamics: this number can be estimated as $n_{pop}\sim ct$. The total cumulative current, then, grows like $J_{cum}\sim n_{pop}J_{ss}=\gamma_{\textsl{g}}ct$; this, in turn, means a larger pumping for the cavity. Once the total pumping rate exceeds the loss rate $\kappa_{c}$, one observes a burst-like increase in $n_{c}$ (right-side plot). This triggers a sudden acceleration of the packet, which quickly moves all the way through the ladder to the final site.
@@
For high values of $\kappa_{c}$ (top plot in Fig.9), we observe a similar, but slightly different process: instead of starting from an empty ladder, we have a non-zero filling, such that the total cavity pumping already exceeds the threshold value $\kappa_{c}$. We observe, therefore, a rapid increase in the cavity population, resulting once again in an acceleration of the ladder dynamics, causing now a dip in the ladder population on the left-hand side. To understand this, consider the dynamics of the leftmost site; in the initial configuration we considered, we have a local equilibrium between the incoming drive current $\gamma_{\textsl{g}}$ and the current $\Gamma(1+\lvert\alpha_{c}\rvert^{2})n_{1}(1+n_{2})$ flowing to the second site. When the cavity population spikes, the latter increase, but not the former, resulting in a population loss on the first site (modes in the center of the ladder, by contrast, are initially unaffected, since both their incoming and outgoing currents are amplified). This initial quench on the first site then quickly affects the neighboring modes, creating the depleted region we observe on the first few modes of the ladder. The decrease in population leads to a lower overall current, making the cavity population decay once more. We are therefore left with a depleted region in the ladder, which is going to propagate at the same effective speed $c$ we discussed above, until it reaches the right edge. Once it starts subsiding, the population moves back towards its initial constant filling configuration, and the current starts increasing again, restarting the cycle.
@@
To summarize, these two opposite regimes of high and low $\kappa_{c}$ exhibit behaviors that are more or less mirror images of one another: in the former, we have a slow creation of a population wave, followed by rapid propagation through the ladder; in the latter, we have a rapid creation of a depletion wave, followed by slow propagation through the ladder. The slow propagation of both kind of waves, however, occurs with the same typical speed $c\sim\sqrt{\Gamma\gamma_{\textsl{g}}}$, which will dominate the overall duration of the process.
The argument given in the main text suggests that the period, rescaled by the effective speed $c\sim\sqrt{\gamma_{\textsl{g}}\Gamma}$, should be compared against the rescaled gain $\gamma_{\textsl{g}}/\kappa_{c}$. In Fig.10, we display the unrescaled and rescaled periods for multiple values of the hopping rate, cavity loss rates, and gain. We observe a very good collapse of all curves, indicating that the ansatz $c\tau=f(\kappa_{c}/\gamma_{\textsl{g}})$ works quite well for both high and low $\kappa_{c}$. However, a more detailed analysis of the function $f$ reveals that it deviates from the behavior $f(\kappa_{c}/\gamma_{\textsl{g}})=\kappa_{c}/\gamma_{\textsl{g}}$ that one would expect from the argument given in the main text. The precise behavior of this function is not fully clear and will be investigated in future works.
Figure 10: Left: unrescaled period versus gain rate for $N=10$ (all frequencies and times are defined with respect to $\kappa_{\ell}$, set here to 11), taking multiple values for the hopping rate and cavity loss: $\Gamma/\kappa_{\ell}=0.5,0.1,0.05$ (blue, green, and red symbols, respectively), and $\kappa_{c}/\kappa_{\ell}=1,5,10,25,50$ (squares, circles, crosses, stars, and diamonds, respectively). Right: rescaled period versus rescaled pumping, with the same parameters. We used this plot for the inset of Fig.3 in the main text.
Appendix D Monte-Carlo simulations
In the absence of any additional Hamiltonian terms, the master equation in Eq. (1) is diagonal in the Fock basis $|\{\vec{n}\}\rangle=|n_{c};n_{1},n_{2}..,n_{N}\rangle$,where $n_{c}$ and $n_{1,..N}$ denote the population of the cavity and the modes in the ladder (i.e., the gain medium), respectively. Therefore, the diagonal elements of the density operator, $P(\{\vec{n}\},t)=\langle\{\vec{n}\}|\hat{\rho}(t)|\{\vec{n}\}\rangle$, which describe the probabilities of different particle configurations, evolve independently from the off-diagonal elements, according to the equation:
@@ Appendix D Monte-Carlo simulations
Where the $dH_{q}$, $dG_{\lambda}$, and $dL_{\lambda}$ are independent random variables, taking binary values $\{0,1\}$, and indicating that a hopping, gain, or loss event, respectively, took place. In a short enough time interval $dt$, the respective probabilities that these variables assume value $1$ are $p(dH_{q}=1)=\Gamma(1+n_{c})n_{q+1}(1+n_{q})dt$, $p(dG_{1}=1)=\gamma_{\textsl{g}}(1+n_{1})dt$, and $p(dL_{\lambda}=1)=\kappa_{\lambda}n_{\lambda}dt$.
The simulation is performed using a standard Gillespie algorithm [9]. By starting from a given initial configuration, $\{n_{p}(t=0)\}$, and evolving a total number of $\mathcal{N}_{t}$ stochastic trajectories in time, we can approximate the expectation value of any function of operators $\hat{n}_{p}$ by an ensemble average. For example,
@@
Appendix E Coherence resonance with other pumping mechanisms
@@
In the main paper, we have mostly focused on a pumping mechanism with $\zeta=1/2$, which corresponds to equal gain and loss on the first ladder mode. Similar effects, however, can also be achieved with other pumping mechanisms. In Fig.11, we show the trajectories, spectra, and coherence factor obtained with $\zeta=1$, which corresponds to pure gain in the first ladder site. The same basic features (presence of bursts in the cavity population, whose coherence peaks for an intermediate value of the gain) can again be observed.